# Look at the Text: Instruction-Tuned Language Models are More Robust Multiple Choice Selectors than You Think

### v1

@@ 1 Introduction
In this paper, we investigate the robustness of the text-based approach. We compare it with the first token answer as well as the debiased first token answer using the first token debiasing method PriDe (Zheng et al., 2023), by measuring the answer robustness under various prompt perturbations. To avoid the labor of feature engineering and the high cost of using a proprietary model as an evaluator, we fine-tune a Mistral-7b-Instruct model and show that it can reliably and accurately detect the choice from the text answers across models and datasets.  [^1] We will share the dataset and trained classifiers at https://github.com/mainlp/MCQ-Robustness.
@@ Acknowledgments
XW, CH and BP are supported by ERC Consolidator Grant DIALECT 101043235 and in parts by Independent Research Fund Denmark (DFF) Sapere Aude grant 9063-00077B. BM is supported by BERD@NFDI (German Research Foundation grant 460037581). PR is a member of the Data and Marketing Insights research unit of the Bocconi Institute for Data Science and Analysis, and is supported by a MUR FARE 2020 initiative under grant agreement Prot. R20YSMBZ8S (INDOMITA).

### v2

@@ 1 Introduction
In this paper, we investigate the robustness of the text-based approach. We compare it with the first token answer as well as the debiased first token answer using the first token debiasing method PriDe (Zheng et al., 2023), by measuring the answer robustness under various prompt perturbations. To avoid the labor of feature engineering and the high cost of using a proprietary model as an evaluator, we fine-tune a Mistral-7b-Instruct model and show that it can reliably and accurately detect the choice from the text answers across models and datasets.  [^1] We share the dataset and trained classifiers at https://github.com/mainlp/MCQ-Robustness.
@@ 3.1 Experimental setup
Test Data Original Options Extra Options Training Data Model Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Mistral Llama2 Gemma* Mistral Llama2 Gemma* AVG MMLU Mistral-7b-Base 0.993 0.870 1.000 1.000 0.997 0.995 1.000 1.000 0.993 0.711 1.000 1.000 0.997 0.929 Mistral-7b-Inst 0.993 0.870 0.997 0.874 0.993 0.992 1.000 1.000 0.997 0.997 0.997 0.997 0.996 0.955 (ID) T5-small 0.986 0.987 0.919 0.913 0.963 0.965 0.986 0.570 0.878 0.665 0.916 0.380 0.941 0.747 T5-base 0.979 0.980 0.997 0.997 0.963 0.964 0.986 0.667 0.878 0.665 0.990 0.664 0.966 0.823 T5-large 0.990 0.870 0.997 0.997 0.980 0.860 0.986 0.570 0.875 0.482 0.983 0.663 0.968 0.740 OpinionQA Mistral-7b-Base 0.993 0.992 0.997 0.997 0.990 0.870 0.979 0.659 0.875 0.537 1.000 1.000 0.972 0.842 Mistral-7b-Inst 0.997 0.873 0.986 0.983 0.977 0.856 0.986 0.665 0.878 0.551 1.000 1.000 0.971 0.821 (OOD) T5-small 0.973 0.972 0.959 0.953 0.786 0.697 0.986 0.795 0.871 0.353 0.920 0.270 0.916 0.673 T5-base 0.986 0.986 0.986 0.863 0.926 0.728 0.979 0.659 0.875 0.395 0.963 0.355 0.953 0.664 T5-large 0.979 0.860 0.993 0.871 0.980 0.768 0.983 0.569 0.864 0.435 0.993 0.796 0.965 0.716
Table 2: Performance of trained text answer classifiers. Both Mistral-7b-Base and Mistral-7b-Inst achieve good performance on test data. *We trained our classifiers only on responses from Mistral-7b and Llama2-7b models while also testing on responses from Gemma-7b.
@@ Acknowledgments
XW, CH and BP are supported by ERC Consolidator Grant DIALECT 101043235 and in parts by Independent Research Fund Denmark (DFF) Sapere Aude grant 9063-00077B. CH is also supported by the DAAD programme Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research. BM is supported by BERD@NFDI (German Research Foundation grant 460037581). PR is a member of the Data and Marketing Insights research unit of the Bocconi Institute for Data Science and Analysis, and is supported by a MUR FARE 2020 initiative under grant agreement Prot. R20YSMBZ8S (INDOMITA).